A 4.5 metre ladder rests against a vertical wall with its feet on level ground 2.5 m from the wall.
(a) How high up the wall does it reach?
(b) If the feet of the ladder are pushed 0.8 m closer to the wall, how much higher up the wall will it reach?

Answer by stanbon(75887)

A 4.5 metre ladder rests against a vertical wall with its feet on level ground 2.5 m from the wall.
------
Draw the picture.
You have a right triangle with base = 2.5 and hypotenuse = 4.5
---------------------
(a) How high up the wall does it reach?
Use Pythagoras:
height^2 + base^2 = hyp^2
h^2 + 2.5^2 = 4.5^2
h^2 = 4.5^2 - 2.5^2
h^2 = 14 ft
h = sqrt(14) ft.
------------------------
(b) If the feet of the ladder are pushed 0.8 m closer to the wall, how much higher up the wall will it reach?
base = 2.5-0.8 = 1.7 ft
hypotenuse is still 4.5
h^2 = 4.5^2 - 1.7^2
h^2 = 17.36
h = sqrt(17.36)
---
height has increased by sqrt(17.36)-sqrt(14) = 0.4249 ft.
===================
Cheers,
Stan H.
===================